OpenStudy (anonymous):

Two cars leave town going in opposite directions. One car is traveling 60 miles per hour (mph), and the other car is traveling 30 mph. How long will it take before the cars are 360 miles apart?

OpenStudy (ganpat):

use the formula speed = distance / time over here speed will be (car1 speed + car 2 speed).. as it is relative and opposite in direction.. therefore, (60 + 30) = 360 / t t = 360/90 = 4 hrs = 240 minutes...